### v1

# A Particle-in-Cell Simulation Framework for Thomson Scattering Analysis in Inertial Confinement Fusion

Abstract
In inertial confinement fusion (ICF), Thomson scattering (TS) is a widely used diagnostic technique for probing plasma conditions. We present a first-principles numerical approach to obtaining scattered light signals of ion acoustic features with high resolution in angle and frequency space using particle-in-cell simulations under typical ICF conditions. Our method demonstrates good agreement with existing theories for thermal collective TS. In the super-thermal collective regime, the results align with theory when the driven plasma modes are well-matched in wave vectors to the probe and collecting beams. Moreover, we also find that TS signals can remain significant even under imperfect wave-vector matching—a result that contradicts the conventional expectation that the TS spectrum strictly follows the plasma density spectrum. We attribute this discrepancy to a beating wave mechanism arising from the interaction between the probe beam and driven plasma density modulations. Our work thus provides a practical framework for interpreting TS signals from driven ion modes, a common yet complex feature in ICF plasmas.
keywords:
particle-in-cell simulation, inertial fusion energy
articletype: Paper
1 INTRODUCTION
In the pursuit of ignition and high-gain inertial confinement fusion (ICF), accurate diagnosis of high-energy-density (HED) plasmas generated by the interaction of intense laser beams with spherical targets or hohlraums is essential. Within these HED plasmas, a variety of physical processes can arise that complicate diagnostics, including the Langdon effect [1, 2, 3], electron thermal transport [4, 5], and laser–plasma instabilities (LPIs)[6, 7, 8, 9, 10, 11, 12, 13]. These processes commonly occur in nonequilibrium or collisional HED plasmas and significantly influence the efficiency of laser energy coupling.
Thomson scattering (TS), the elastic scattering of low-energy photons by electrons, has become a well-established diagnostic technique for probing HED plasma conditions, driven waves and distribution functions in ICF. Collective Thomson scattering (CTS) arises when the scattered waves from individual electrons interfere coherently, allowing access to information about the collective behavior of the plasma.[14]
The power of the scattered light of CTS, denoted as $P_{s}(\vec{R},\omega_{s})$, reflects characteristics of electron density perturbations. The scattered power in the unit frequency $d\omega_{s}$ and unit solid angle $d\Omega$ is given by[14]:
$P_{s}(\vec{R},\omega_{s})d\Omega d\omega_{s}=\frac{P_{i}e_{0}^{2}}{A2\pi}d\Omega d\omega_{s}\left(1+\frac{2\omega}{\omega_{i}}\right)NS(\vec{k},\omega),$ (1)
where $P_{i}$ denotes the power of the incident (probe) beam, $r_{0}$ is the classical electron radius, and $A$ is the cross-sectional area of the scattering volume. $N$ represents the number of electrons within this volume. $\omega_{i}$ and $\omega_{s}$ are the frequencies of the incident and scattered light, respectively. $\omega$ and $\vec{k}$ denote the frequency and wave vector of the electron density fluctuations modes that satisfy the wave–wave matching conditions:
$\vec{k}_{s}$ $=\vec{k}_{i}\pm\vec{k},$ (2)
$\omega_{s}$ $=\omega_{i}\pm\omega,$
The dynamic form factor $S(\vec{k},\omega)$ is defined by:
$S(\vec{k},\omega)=\lim_{V\rightarrow\infty,T\rightarrow\infty}\frac{1}{VT}\bigg<\frac{|n_{e}(\vec{k},\omega)|^{2}}{n_{e0}}\bigg$>, (3)
where $\langle\cdot\rangle$ denotes an ensemble average, $n_{e0}$ represents the unperturbed average electron density, $n_{e}(\vec{k},\omega)$ is the electron density fluctuation spectrum, $V$ and $T$ are the spatial volume and temporal duration of the measurement, respectively.
Thermal CTS is one subtype of CTS and an essential diagnostic tool, corresponding to the scattering process driven by thermal fluctuations. According to the fluctuation-dissipation theorem, the corresponding spectrum in a collisionless thermal equilibrium plasma takes the form[14]:
$S(\vec{k},\omega)=\frac{2\pi}{k}\bigg|1-\frac{\chi_{e}}{\epsilon}\bigg|^{2}f_{e}\bigg(\frac{\omega}{k}\bigg)+\frac{2\pi Z}{k}\bigg|\frac{\chi_{e}}{\epsilon}\bigg|^{2}f_{i}\bigg(\frac{\omega}{k}\bigg),$ (4)
Where $\chi_{e}$ and $\chi_{i}$ are the susceptibilities of electron and ion respectively, $\epsilon=1+\chi_{e}+\chi_{i}$ is the plasma dielectric function, $f_{e}$ and $f_{i}$ represent the electron and ion velocity distribution function respectively. These distributions embody key plasma parameters, including density, temperature, and flow velocity. Consequently, by fitting measured thermal CTS spectra, one can infer these plasma properties[15, 16, 17, 18, 19, 20, 21, 22, 23].
However, HED plasmas under realistic ICF conditions often deviate from the collisionless thermal equilibrium assumption. Examples include high-Z collisional plasmas[24], plasmas with heat flux[5], and those exhibiting super-Gaussian electron distributions[25, 26]. In such regimes, conventional CTS theories become inaccurate, complicating experimental data analysis and diagnostic interpretations. This calls for new approaches to analyze how CTS spectra depend on these conditions.
Particle-in-cell (PIC) simulation is a first-principles method that self-consistently solves Maxwell’s equations and tracks particle motion under electromagnetic forces [27], making it well-suited for studying kinetic effects in nonthermal or collisional plasmas. PIC simulations have been used to examine non-collective TS[28]. For CTS, pioneer investigations have proposed 1D and 2D PIC methods to study nonthermal plasmas [29] along certain directions. While recent experiments seeking to obtain detailed plasma properties or to diagnose non-Maxwellian plasmas via collective Thomson scattering (CTS) often rely on collecting scattered light from multiple angles [30, 31], high-resolution, angle- and frequency-resolved data are still required to guide experimental design. To integrate experimental TS measurements, development of numerical simulation methods with sufficient angular and spectral resolution is essential.
Another application for CTS in ICF is the diagnosing of LPI, a key process that strongly relates with driven plasma waves. For example, in indirect-drive ICF, ion density modulations induced by the beating of multiple overlapping laser beams at the laser entrance hole of the hohlraum can significantly affect cross-beam energy transfer—a critical mechanism for implosion performance in indirect drive ICF[6]. Similarly, perturbations driven by the Langmuir decay instability play a pivotal role in the growth and saturation of two-plasmon-decay instability, influencing hot‑electron generation[32, 33, 34]. The CTS measuring these externally driven plasma waves is called as super-thermal collective Thomson scattering (SCTS). Because the amplitude of driven perturbations significantly exceeds that of thermal fluctuations, SCTS signals are typically orders of magnitude stronger than thermal CTS. To support experimental studies of these processes, it is therefore essential to develop numerical methods that properly resolve angular variations in the SCTC signal under such conditions.
In this work, we developed the method to obtain CTS light spectra with high angular and spectral resolution from 2D PIC simulations, enabling the detailed analyze of nonequilibrium anisotropic plasmas in future studies. The ion-acoustic features in the simulated spectra are clearly resolved and exhibit shapes that agree well with theoretical predictions. We further perform a systematic comparison between SCTS signals and the underlying driven density perturbations, revealing that perturbations with wave vectors mismatched from the scattering geometry still yield significant TS signals—an effect not captured by standard theory. Our analysis suggests this should result from a beating-wave interaction between the probe beam and the driven density perturbation.
Figure 1: Scheme of PIC simulation setup used to obtain TS signal, along with the wave vector of probe beam $\vec{k}_{i}$, the scattered light $\vec{k}_{s}$ at scattering angle $\theta_{s}$ and the matched density perturbation $\vec{k}$.
The remainder of this paper is organized as follows: Section II describes the numerical method of obtaining CTS signals with high resolution in angular and frequency space; Section III presents the capability of our method, including numerical results of CTS spectra with different level of ion-ion collisional effects; Section IV focuses on SCTS, in which the new phenomenon, drawing attention to a newly observed phenomenon in which scattered light results from perturbations with mismatched wave vectors; Section V summarizes the paper and outlines directions for future work on studying ICF plasmas using TS diagnostics.
2 SIMULATION PROCEDURE TO OBTAIN THOMSON SCATTERING SIGNAL
In this section, we demonstrate the general procedure to obtain CTS signals with high frequency and angular resolutions using PIC simulation. The probe beam and plasma properties, along with temporal and spatial grid setting and sampling setting for later Fourier analysis to obtain TS signal are described. PIC simulations in this article are conducted using OSIRIS code[35]. An example to obtain CTS spectra in an equilibrium uniform plasma is presented.
2.1 PIC simulation setup
The simulation is performed in two dimensions (the $xy$-plane) on a domain ranging from $200\,c/\omega_{0}\times 200\,c/\omega_{0}$ ($11.2\,\mu m\times 11.2\,\mu m$) to $800\,c/\omega_{0}\times 800\,c/\omega_{0}$ ($44.7\,\mu m\times 44.7\,\mu m$), where $c$ denotes the speed of light and $\omega_{0}$ denotes the frequency of the laser beam with a wavelength of 351 nm. The domain is uniformly filled with hydrogen plasma at electron density $n_{e}=0.05\,n_{c}$, where $n_{c}$ is the critical density for the 351 nm laser. The initial electron and ion temperatures are $T_{e}=2\,keV$ and $T_{i}=0.5\,keV$, respectively.
A probe beam is injected from the left boundary and propagates along the x-direction, as illustrated in Fig.1. It has a frequency of $\omega_{i}=\tfrac{4}{3}\omega_{0}$ (corresponding to a 263 nm wavelength) and is linearly polarized perpendicular to the xy-plane. The transverse profile of the beam is Gaussian, with a peak intensity $I=9\times 10^{12}\,\mathrm{W/cm^{2}}$ and a full width at half maximum (FWHM) varying from $100\,c/\omega_{0}$ ($5.6\,\mu m$) to $400\,c/\omega_{0}$ ($22.3\,\mu m$).
The spatial grid resolution is $\Delta x=0.25\,c/\omega_{0}$ ($0.014\,\mu m$), and the temporal step size is $\Delta t=0.17\,\omega_{0}^{-1}$ ($0.003\,\text{fs}$). Each grid cell contains 100 macroparticles representing electrons and another 100 for ions. Particle boundary conditions are thermal, while electromagnetic fields satisfy open boundary conditions.
2.2 Frequency- and angular-resolved TS spectrum
To obtain the scattered-light spectrum $P_{s}(\theta,\omega)$ from PIC simulations, the $E_{z}$ field component is extracted. But the raw output $E_{z}(x,y,t)$ includes abundant data that can complicate analysis. To address this, OSIRIS’s time‑FFT module is applied[36, 37], which performs in-situ Fast Fourier Transforms during the simulation and outputs only the targeted frequency range, yielding $E_{z}(x,y,\omega)$ and substantially reducing data volume. Provided the required resolutions in $k$- and $\omega$-space are met, sampling in the x, y and t domain further reduces the data volume without compromising the spectral analysis.
Figure 2: Thomson scattering spectra for scattering angles of $\theta_{s}$ of (a) $60^{\circ}$, (b) $90^{\circ}$, (c) $120^{\circ}$, and (d) $150^{\circ}$, each averaged over $\pm 1^{\circ}$.
The spatial sampling interval is $4\Delta x$, and the temporal sampling interval is $2\Delta t$, corresponding to a measure range of $3.14,3.14]\,\omega_{0}/c$ for $k$ and $9.24,9.24]\,\omega_{0}$ for $\omega$. For the Thomson-scattering spectrum exhibiting ion-acoustic features generated by the $\tfrac{4}{3}\omega_{0}$ probe beam, we select the frequency window $1.32\,\omega_{0},\,1.35\,\omega_{0}]$, which captures the relevant spectral signatures. The simulation duration is set to $175{,}000\,c/\omega_{0}$ (approximately $50\text{ ps}$), enabling a frequency resolution of $d\omega=2.3\times 10^{-5}\,\omega_{0}$. This value is significantly finer than the resolution typically achieved in experiments[19], where ion-acoustic features in CTS spectra are measured with a spectral resolution of about $0.06nm$ , corresponding to $3.0\times 10^{-4}\omega_{0}$. The wavenumber resolution is $dk=7.8\times 10^{-3}\omega_{0}/c$ for a simulation domain size of $800\,c/\omega_{0}\times 800\,c/\omega_{0}$ ($44.7\mu m\times 44.7\mu m$).
Figure 3: (a) Scattered spectra at $\theta_{s}=90^{\circ}$ from simulations with different random seed. (b) Angular distribution of height differences from averaged results of 1, 4, 16 simulations. (c) Height difference $\overline{|\Delta h|}$ decay with the number of averaging simulations $N_{simu}$ in a manner of $1/\sqrt{N_{simu}}$. (d) Scattered spectra at $\theta_{s}=90^{\circ}$, averaged over 16 simulations.
Though a spatial FFT, $E_{z}(x,y,\omega)$ output is transformed into $E_{z}(k_{x},k_{y},\omega)$ data. Next, a linear interpolation converts this to polar coordinates: $E_{z}(k,\theta,\omega)$. By applying the dispersion relation for light, $\omega^{2}=k^{2}c^{2}+\omega_{pe}^{2}$, where $\omega_{pe}$ is the plasma frequency, we obtain the angular–spectral distribution $E_{z}(\theta,\omega)$ for scattered light.
3 ION-ACOUSTIC FEATURES OF THERMAL THOMSON SCATTERING OBTAINED FROM SIMULATIONS
We apply our method to thermal CTS cases. We start from the equilibrium collision-less case, where TS spectrum can be conveniently described with existing theories.
3.1 Ion-acoustic features of Thomson scattering
For the plasma conditions described in Section II, the frequency of ion-acoustic peaks $\omega_{s}$ in the TS spectrum can be calculated using:
$\omega_{s}=\omega_{i}\pm\omega_{iaw}=\omega_{i}\pm kc_{s},$ (5)
where $c_{s}\approx\sqrt{(ZT_{e}+3T_{i})/m_{i}}\approx 1.8\times 10^{-3}c$ is the ion sound speed in this simulation, $k=k_{s}-k_{i}\approx 2k_{i}\sin(\theta_{s}/2)$ is the wave number of the fluctuation matching with the scattering angle $\theta_{s}$.
Following the method mentioned in Section II.C, we obtain low-frequency part of Thomson scattering spectra at different scattering angles of $60^{\circ}\pm 1^{\circ}$, $90^{\circ}\pm 1^{\circ}$, $120^{\circ}\pm 1^{\circ}$, and $150^{\circ}\pm 1^{\circ}$, as shown in Fig.2. To reduce spectral noise, second-order smoothing with a window size of $3.0\times 10^{-4}\omega_{0}$ is applied iteratively to these spectra until their shapes converge. The simulated TS spectra are in good agreement with the theoretical profiles calculated from Eq. (4), which indicates that our method can produce valid TS spectra with high angular and spectral resolution.
3.2 Achieving higher signal-to-noise ratio
Although the method described above appears straightforward, the signal-to-noise ratio presents a primary challenge in its practical application. This issue arises from the limited statistical sampling in PIC simulations. Under realistic experimental conditions, CTS spectra represent a statistical average over a three-dimensional volume of hundreds of micrometers. In contrast, PIC simulations typically operate over two-dimensional domains of only tens of micrometers. This disparity introduces much stronger fluctuations into the simulated spectra, which can significantly affect the signal-to-noise ratio of the obtained CTS spectra.
One clear example is the variation in peak heights of the ion-acoustic features, which are sensitive to the drift velocity between electrons and ions. In an isotropic plasma, the red-shifted and blue-shifted ion-acoustic peaks should be symmetric. However, in a simulation with a $200\,c/\omega_{0}\times 200\,c/\omega_{0}$ ($11.2\,\mu m\times 11.2\,\mu m$) domain, the peak height asymmetry can be significant, as shown in Fig.3(a).
In these simulations, the initial thermal velocities of particles are randomly sampled from a Maxwellian distribution using an initial random seed. Changing this random seed does not alter macroscopic plasma quantities, such as density and temperature, but it does affect the microscopic realizations of the system. As shown in Fig. 3(a), the pattern of height differences varies with the random seeds used to sample the initial thermal velocities, indicating that CTS spectra are affected by the initial microscopic realizations—a result that can be attributed to limited statistical sampling. We quantify this asymmetry using the normalized height difference:
$\Delta h=\frac{h_{R}-h_{B}}{h_{R}+h_{B}}$ (6)
where $\Delta h$ is the normalized difference, and $h_{R}$ and $h_{B}$ are the heights of the red-shifted and blue-shifted peaks, respectively. Figure 3(b) includes the angular distribution of $\Delta h$ from a single simulation, where $\Delta h$ varies in different angles randomly. This randomness can obscure the measurement of drift velocities, thereby reducing the reliability of physical diagnostics, such as those involving electron heat transport.
In Sec II.A, we demonstrated that simulations with a domain size of $800\,c/\omega_{0}\times 800\,c/\omega_{0}$ ($44.7\,\mu m\times 44.7\,\mu m$) is adequate to yield ion-acoustic peaks of nearly equal height and a significantly improved signal-to-noise ratio. But the required simulation space may vary when different physical subjects are studied. So a more general approach to improve statistical robustness is required. This is accomplished by conducting multiple simulations under identical macroscopic conditions but with varied microscopic realizations (e.g., different random seeds for initial velocity sampling), and then averaging their resulting spectra.
Figure 4: Simulated Thomson scattering spectrum with different ion-ion collision rate.
The angular distribution of $\Delta h$ for different numbers of averaged simulation results, $N_{simu}$, is shown in Fig. 3(b), where $\Delta h$ decreases significantly as $N_{simu}$ increases. A more quantitative result is illustrated in Fig.3(c), in which the average height difference $\overline{|\Delta h|}$ decreases following a $1/\sqrt{N_{simu}}$ scaling law. The spectrum obtained by averaging 16 simulations is shown in Fig.3(d), where the noise level is significantly reduced compared to Fig. 3(a).
3.3 Capability to simulate ion-ion collision’s influence on Thomson scattering spectrum
Ion–ion collisions modify the ion-acoustic features of the TS spectrum by enhancing collisional damping and reducing Landau damping [24]. Their influence on TS spectrum with wave vector $k$ can be evaluated with the parameter $k\lambda_{ii}$. Here $\lambda_{ii}$ is the ion mean free path due to ion-ion collisions given by $\lambda_{ii}=v_{ti}/\nu_{ii}$, where $v_{ti}$ represents ion thermal velocity and $\nu_{ii}$ denotes ion-ion collision frequency. The condition $k\lambda_{ii}\gg 1$ corresponds to the collisionless cases, $k\lambda_{ii}\sim 1$ to the intermediate cases, and $k\lambda_{ii}\ll 1$ to the strongly collisional cases.
In PIC simulations, we can adjust the level of collision by setting the Coulomb logarithm manually[38, 39]. A series of simulations is conducted with collision parameters $k\lambda_{ii}=0.1,1,10,\infty$. Here $k\lambda_{ii}=\infty$ correspond to the case where ion-ion collision is turned off in PIC simulation. Simulation domain with a size of $400\,c/\omega_{0}\times 400\,c/\omega_{0}$ ($22.3\,\mu m\times 22.3\,\mu m$) is filled with uniform hydrogen plasma. Electron and ion temperatures are $T_{e}=2\,keV$ and $T_{i}=1\,keV$, respectively. Rest of parameters are the same with previous context. Simulated spectra with scattering angle of $90^{\circ}\pm 1^{\circ}$ is shown in Fig.4.
In the collisionless case, the ion-acoustic waves undergo strong Landau damping due to $ZT_{e}/T_{i}=2$ in these simulation, rendering the resonance peaks for $k\lambda_{ii}\sim\infty$ and $k\lambda_{ii}=10$ insignificant. When $k\lambda_{ii}=1$ and $k\lambda_{ii}=0.1$, strong collisionality suppresses ion Landau damping, and the resonance peaks become pronounced. In the strongly collisional regime, an additional resonance peak at zero frequency shift, corresponding to an entropy wave, appears. These tendencies in the simulated spectra are consistent with the theoretical predictions [24], confirming that the simulation properly reproduces the effect of collisions.
4 SUPER-THERMAL THOMSON SCATTERING VIA PIC SIMULATION
Unlike thermal CTS, SCTS arises from light scattering off driven plasma waves whose amplitudes far exceed thermal fluctuations, making it less susceptible to signal-to-noise limitations in PIC simulations. In such simulations, the amplitude, wave vector, and frequency of the driven waves can be precisely controlled, enabling reproduction of the SCTS process and facilitating a targeted study of the matching condition in Eq. (2).
Figure 5: (a) Two laser beams beating to drive an ion density perturbation. (b) TS spectrum at $60^{\circ}$ with SCTS peak corresponding to the driven perturbation frequency. (c) TS spectrum of three perturbations with different frequencies but the same wave vector. (d) Relation between the SCTS power and the amplitude of the corresponding perturbation.
In this section, we show that when the matching condition is well satisfied, the frequency distribution of the simulated SCTS signal reflects the frequency of the driven perturbation, while its angular distribution reflects the perturbation’s wave vector. We also identify a new phenomenon in which scattered light originates from density perturbations with a mismatched wave vector.
4.1 Super-thermal Thomson scattering from matched density perturbations with different frequency
To investigate SCTS induced by matched density perturbations, we generate these perturbations through laser beating. When two lasers $(\vec{k}_{drive,1},\omega_{drive,1})$ and $(\vec{k}_{drive,2},\omega_{drive,2})$ overlap spatially, their interference envelope has a wave vector and frequency of $(\vec{k}_{drive,1}-\vec{k}_{drive,2},\omega_{drive,1}-\omega_{drive,2})$. The ponderomotive force associated with this interference pattern drives a density fluctuation. By adjusting the wave vectors and frequencies of the beams, density perturbations with arbitrary wave vector and frequency can be generated.
The simulations are performed in a domain of $200\,c/\omega_{0}\times 200\,c/\omega_{0}$ ($11.2\,\mu m\times 11.2\,\mu m$), filled with uniform hydrogen plasma. The plasma and probe parameters are the same as those described in Sec.II. As shown in Fig.5(a), two laser beams, each with an intensity of $3.5\times 10^{14}\,\mathrm{W/cm^{2}}$ and a wave vector of $0.93\,\omega_{0}/c$, overlap at an angle of $90^{\circ}$, driving an IAW with a wave vector $k_{\mathrm{iaw}}$ matched to $60^{\circ}$ Thomson scattering. The frequency difference between the two driving beams corresponds to the IAW eigenfrequency, $\omega_{\mathrm{iaw}}=k_{\mathrm{iaw}}c_{s}$. The corresponding SCTS signal at $60^{\circ}$ is shown in Fig. 5(b), where a narrow-band peak appears at the matched frequency $\omega_{i}+\omega_{\mathrm{iaw}}$.
We further excite three ion density perturbations with the same wave vector but different frequencies in a single simulation. The configuration is the same as in Fig.5(a), except that one drive beam $(\vec{k}_{\mathrm{drive},1},\omega_{\mathrm{drive}}+\omega_{\mathrm{iaw}})$ is replaced by three overlapping beams with frequencies $\omega_{\mathrm{drive}}-\tfrac{1}{2}\omega_{\mathrm{iaw}}$, $\omega_{\mathrm{drive}}$, and $\omega_{\mathrm{drive}}+\tfrac{1}{2}\omega_{\mathrm{iaw}}$, respectively. The intensity of each drive beam is set to $1\times 10^{14}\,\mathrm{W/cm^{2}}$. Three SCTS signals are simultaneously observed in the $60^{\circ}$ scattering spectrum, with frequency shifts corresponding to the driven perturbations, as shown in Fig.5(c).
Figure 6: (a) Initial electron density distribution with four density perturbations at the left half of simulation domain. (b) Angular distribution of zero-frequency-shifted signal. The blue solid line is the distribution in simulation result, the gray dashed lines are the theoretical position of SCTS peaks.
A series of simulations is performed using the same configuration described above, but with varying drive-beam intensities and frequency differences. These variations generate density perturbations of different amplitudes, leading to corresponding changes in the SCTS signals. By comparing the amplitudes of the SCTS signals with the perturbation amplitudes, we find that the scattered power is proportional to the square of the perturbation amplitude, as shown in Fig. 5(d). This result indicates that the frequency distribution of the TS signal directly reflects the frequency distribution of the density perturbations when the matching condition is satisfied.
4.2 Super-thermal Thomson scattering from matched density perturbations with different wave vector
When driving density perturbations through laser beating, the laser overlap region occupies only a small fraction of the total simulation area, leading to inefficient use of computational resources. An alternative method of generating electron density fluctuations is initializing them at the start of the simulation without injecting ions in the simulation domain. Owing to the nature of PIC simulations, which automatically enforce the Poisson equation [27], this initialized electron density distribution is sustained throughout the simulation. This approach can produce stationary perturbations effectively.
To investigate SCTS arising from matched density perturbations with different wave vectors, we initialize four electron density fluctuations within a single simulation, as shown in Fig. 6(a). The simulation is conducted in a domain of $400\,c/\omega_{0}\times 1800\,c/\omega_{0}$ ($22.4\,\mu\mathrm{m}\times 100.6\,\mu\mathrm{m}$), filled with a background electron density of $n_{0}=0.05n_{c}$. The boundaries perpendicular to the $y$-axis are set to be periodic, while the remaining boundaries were thermal for particles and open for electromagnetic fields.
The density perturbations are initialized in the left half of the simulation domain (see Fig. 6(a)). These perturbations have wavelengths of $20\,\mu\mathrm{m}$, $10\,\mu\mathrm{m}$, $5\,\mu\mathrm{m}$ and $2\,\mu\mathrm{m}$, each oriented such that its wave vector matched that of Thomson-scattered light at a different scattering angle. Their amplitudes are set to $(1,0.5,0.25,0.1)\times 10^{-3}n_{0}$, respectively.
A plane wave with frequency $\omega_{i}=\tfrac{4}{3}\omega_{0}$ is launched from the left boundary along the $x$-axis as a probe beam. The scattered signal is collected in the right half of the simulation domain. This configuration, which separates the interaction region from the detection region, is designed to distinguish far-field electromagnetic signals from near-field ones and will also be discussed in the next section. The angular distribution of the TS signal is shown in Fig. 6(b), where SCTS features are observed at different angles that agree with those predicted by the matching condition. This result demonstrates that the TS spectrum effectively represents the wave vector distribution of matched electron density fluctuations.
4.3 Thomson scattering from perturbations with a wave vector deviated from matching condition
The SCTS under matching conditions has been presented in the previous sections. When measuring driven plasma waves using SCTS, the scattering geometry should be carefully designed to satisfy the matching condition. However, plasma conditions and the resulting driven plasma waves are often difficult to predict, and the matching conditions may not always be fulfilled in experiments. To investigate a representative scenario relevant to ICF plasmas, where Thomson scattering is employed to diagnose laser-beat–driven ion waves, we perform a PIC simulation to examine this process under a mismatched condition. In the simulation shown in Fig. 7(a), two drive beams with intensity of $10^{14}W/cm^{2}$ and frequency of $\omega_{0}$ is used to generate density perturbations through beating. A probe beam, with a frequency $\omega_{s}=\frac{4}{3}\omega_{0}$ and intensity of $10^{13}W/cm^{2}$, is launched from the lower boundary at $40.5^{\circ}$ clockwise from the y-axis, passing through the region where drive beams overlap. The scattered light is collected along y-axis at scattering angle $\theta_{s}=40.5^{\circ}$. The driven density perturbation has a wave vector $\vec{k}_{beat}$ that slightly differs from the matched wave vector corresponding to the measured scattering angle, $\vec{k}_{match}$. The deviation is quantified as $\vec{k}_{beat}-\vec{k}_{match}=0.08\,\vec{k}_{s}$. The corresponding density perturbation distribution displayed in Fig. 7(b) shows no signal at the matched position. However, in the scattering spectrum shown in Fig. 7(c), a super-thermal Thomson scattering signal with an amplitude significantly higher than thermal TS signal is observed. This result indicates that the observable scattered light reflects not only the amplitude of the matched mode but also contributions from surrounding unmatched modes. This introduces additional challenges in interpreting SCTS signals from driven ion waves.
Figure 7: (a) Distribution of electric field $E_{z}$ in the simulation domain, along with the wave vectors of drive beams ($\vec{k}_{drive,1}$ and $\vec{k}_{drive,2}$), the driven density perturbation ($\vec{k}_{beat}$), the probe ($\vec{k}_{i}$), the collected scatter light ($\vec{k}_{s}$) and the matched density perturbation ($\vec{k}_{match}$). (b) Electron density perturbation in the phase space $(k_{x},k_{y})$. (c) SCTS spectrum along the y-direction ($\theta_{s}=40.5^{\circ}$), the position of signal corresponding to $\vec{k}_{beat}$ is marked by dashed line.
Figure 8: (a) Spatial distribution for the envelope of perturbed current density $\vec{J}_{\rm pert}$, along with the distribution of excited electromagnetic field profiles for (a) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=\omega_{0}/c$ and (b) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=0.5\,\omega_{0}/c$.
We provide an explanation for the SCTS emission from unmatched ion density perturbations. We consider the interaction of the probe beam electric field $\vec{E}_{i}\sin(\vec{k}_{i}\cdot\vec{x}-\omega_{i}t)$ with electron density perturbation $\delta n_{e}\sin(\vec{k}_{0}\cdot\vec{x}-\omega_{0}t)$. A current density perturbation can be directly generated via wave beating: $\vec{J}_{pert}=\frac{\vec{E}_{i}e\delta n_{e}}{2m_{e}}(\cos[(\vec{k}_{i}-\vec{k}_{0})\cdot\vec{x}-(\omega_{i}-\omega_{0})t]-\cos[(\vec{k}_{i}+\vec{k}_{0})\cdot\vec{x}-(\omega_{i}+\omega_{0})t])$. This $\vec{J}_{pert}$, with wave vector $\vec{k}_{pert}=\vec{k}_{i}\pm\vec{k}_{0}$ and frequency $\omega_{pert}=\omega_{i}\pm\omega_{0}$, could excite electromagnetic waves[40]. Starting from the Maxwell equation, we obtain the relation between currents and electric fields:
$\nabla\times(\nabla\times\vec{E})+\frac{1}{c^{2}}\frac{\partial^{2}\vec{E}}{\partial t^{2}}=-\frac{4\pi}{c^{2}}\frac{\partial\vec{J}}{\partial t},$ (7)
in which $\vec{J}$ includes $\vec{J}_{pert}$, the perturbation current driven by the interaction between the probe beam and the density perturbations, and $\vec{J}_{cond}$, the conduction current arising from the interaction of background electrons with the electric field. For high frequency electromagnetic wave in a non-magnetized plasma, $\vec{J}_{cond}$ satisfies $\frac{\partial\vec{J}_{cond}}{\partial t}=\frac{e^{2}n_{e}}{m_{e}}\vec{E}$. Substitute $\vec{J}=\vec{J}_{pert}+\vec{J}_{cond}$ into Equation (9), we get:
$\nabla\times(\nabla\times\vec{E})+\frac{1}{c^{2}}\frac{\partial^{2}\vec{E}}{\partial t^{2}}=-\frac{\omega_{pe}^{2}}{c^{2}}\vec{E}-\frac{4\pi}{c^{2}}\frac{\partial\vec{J}_{pert}}{\partial t}.$ (8)
Then we have:
$(\nabla^{2}-\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}-\frac{\omega_{pe}^{2}}{c^{2}})\vec{E}=\frac{4\pi}{c^{2}}\frac{\partial\vec{J}_{pert}}{\partial t}.$ (9)
Therefore, the scattered light associated with unmatched ion density perturbations in a plasma can be approximated as the radiation of $\vec{J}_{\mathrm{pert}}$ in vacuum. We insert the following external electron current into a fully vacuum one-dimensional PIC simulation:
$\vec{J}_{\rm{pert}}=\vec{J}_{0}\sin(k_{\rm{pert}}x-\omega_{\rm{pert}}t)\exp\left[-(\frac{x-x_{0}}{L_{\rm{pert}}})^{6}\right],$ (10)
where $L_{\rm pert}=700\,c/\omega_{0}$ is the length of the perturbation region, and $x_{0}=980\,c/\omega_{0}$ represents its center in the simulation domain. We set the current to zero beyond the width $L_{pert}$. Figures 8 show the spatial envelope of $\vec{J}{\rm pert}$ along with the resulting electromagnetic field for two cases: matched, with $\omega_{\rm pert}/k_{\rm pert}=c$, and mismatched, with $\omega_{\rm pert}/k_{\rm pert}=2c$.
In the matched case, a propagating electromagnetic wave is excited within the perturbed region and radiates outward. By contrast, the mismatched case features a near-field electromagnetic perturbation localized to the driven region; at the boundary, scattered light at frequency $\omega_{pert}$ is emitted into the surrounding vacuum. These results indicate that scattered light can still be generated even under unmatched wave-vector.
Since driven perturbations that deviate from the matching condition can still produce SCTS signals, the actual ion modes underlying the diagnosed SCTS signals should not be restricted to the exact matching condition. This needs to be carefully investigated and taken care of in analyzing our experimental results on Shenguang facility.
5 Summary
In summary, we have developed a first-principles numerical approach for obtaining scattered light signals of ion acoustic features with high angular and frequency resolution under typical ICF conditions, using PIC simulations. The method reproduces existing theoretical predictions for thermal collective Thomson scattering, and extends to the super-thermal regime where driven plasma modes are present. In cases where the driven modes are well-matched in wave vectors with the probe and collection geometries, the results agree well with theoretical expectations. Importantly, we find that significant TS signals can persist even under imperfect matching conditions, in contrast to the conventional knowledge that the TS spectrum strictly follows the plasma density spectrum. This discrepancy is explained by a beating-wave mechanism arising from the interaction between the probe beam and the driven plasma density modulations. These findings highlight that SCTS signals can carry contributions from both matched and unmatched modes, and thus provide a practical framework for interpreting scattering spectra in the presence of driven ion modes, which are a common yet complex feature in ICF plasmas.
In this work, our analysis has focused on the ion-acoustic features of the Thomson scattering spectrum, which are typically the dominant components under ICF conditions. However, the electron feature is also an important diagnostic target in experiments. Because its intensity is several orders of magnitude lower than that of the ion-acoustic feature, accurately resolving it in PIC simulations remains challenging due to statistical noise. In future work, we plan to employ our simulation method to obtain electron features. We also aim to apply the present framework to more specific ICF-relevant problems, such as diagnosing CBET-related density perturbations in laser entrance hole region of hohlraums using SCTS.
[…]
This work was supported by the National Key R&D Program of China, No. 2023YFA1608400 and National Natural Science Foundation of China (NSFC) (Grant No. 12275269, U2430207, 12275251). The numerical simulations in this paper were conducted on Hefei advanced computing center.